A sample of a gas occupies a volume of 29.2 L at 11ºC and 793 torr. What will be the new volume if the temperature increases to 71ºC and the pressure decreases to 434 torr? New volume: L
Answer #1

Given:

Pi = 793 torr

Pf = 434 torr

Vi = 29.2 L

Ti = 11.0 oC

= (11.0+273) K

= 284 K

Tf = 71.0 oC

= (71.0+273) K

= 344 K

use:

(Pi*Vi)/(Ti) = (Pf*Vf)/(Tf)

(793 torr*29.2 L)/(284.0 K) = (434 torr*Vf)/(344.0 K)

Vf = 64.6 L

Answer: 64.6 L